Express the following linear equations in the form ax + by + c = 0 and indicate the values of a, b and c in each case:

 

 -2x+3y = 6

 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

–2x + 3y = 6

Re-arranging the given equation, we get,

–2x + 3y – 6 = 0

The equation –2x + 3y – 6 = 0 can be written as,

(–2)x + 3y +(– 6) = 0

Now comparing (–2)x + 3y +(– 6) = 0 with ax + by + c = 0

We get, a = –2

b = 3

c = -6

Thanks